### v2

@@ 2.3 Experimental results
The alkali cell is pumped by a sequence of pump pulses from a grating-stabilized diode laser. Inside the magnetic shield, a set of three coils provides a rotating magnetic field and a leading magnetic field. The cell is heated by an electric heater driven by AC at a frequency of 131.5 kHz. The AC heater is turned on during the pump time and turned off during the measurement time to reduce magnetic noise from the heater.
As shown in Fig. 1a, a rotating magnetic field with an amplitude of approximately 18 $\mathrm{\mu T}$ is applied in the transverse plane, and a leading magnetic field $B_{z}$ is applied along the longitudinal direction. The total amplitude of the applied magnetic field is maintained at approximately 50 $\mathrm{\mu T}$. The frequency of the pump beam pulse is set to 348 kHz, synchronized with the Larmor frequency to achieve the highest pumping rate. The duty cycle of a pump beam pulse is approximately 1.4%.
After the pump time, the pump beam is turned off, and the free induction decay (FID) signal of the spin precession is measured by a linearly polarized probe beam originating from a vertical cavity surface emitting laser (VCSEL). The optical rotation of the linearly polarized probe beam is measured by a balanced polarimeter consisting of a Wollaston prism and a quadrant photodiode. The signal from the photodiode is then passed through a differential low-noise amplifier and a high-pass filter with a cutoff frequency of 150 kHz.
@@
2.3.1 Results of the systematics
[Probe beam heading error] To study the probe beam heading error, the sensor rotates about the y-axis in the x-z plane with an angle of $\alpha$. This error can induce a magnetic field measurement systematic that depends on the angle $\alpha$, and its sign depends on the rotating direction of the rotating field. The measured probe heading error when $\omega_{m}=2\pi\times 480$ Hz is shown in Fig. 2(a). The theory is based on Eq. 7. “Theory 1” is for anti-clockwise rotations (Shot 1 and Shot 3 in Table 2), which is $-\omega_{m}/\gamma\cdot\tan\alpha$. “Theory 2” is for clockwise rotations (Shots 2 and 4 in Table 2), which is $\omega_{m}/\gamma\cdot\tan\alpha$. The probe beam heading error can be canceled out by averaging two opposite rotating field measurements.
[Berry’s phase] When $\theta\approx 21^{\circ}$, magnetic field measurement systematic caused by Berry’s phase shift is measured to be 5.1 nT, while Eq. 5 predicts 4.7 nT, as shown in Table. 3.
[Second harmonic] The amplitude of the second harmonic of $\omega_{m}$ in phase shift is 35 mrad, which matches Eq. 6 prediction well.
[Static heading error] The static heading error is measured when the magnetometer sensor rotates an angle $\beta$ = 24 in the y-z plane which leads to a heading error $B_{SH}=\mathcal{B}_{H}\sin\beta$, we set nuclear Landé factor $g_{I}\approx 0$ [38] and
where $P$ is spin polarization, $A_{hf}=h\cdot 3.417$ GHz is the hyperfine structure constant for ground state [39], $h$ is Planck constant. We can estimate a static heading error of 2.7 nT based on these equations (P=78%, the initial polarization is estimated using the decay feature of the FID signal envelope [40]), the experimental result is 2.7 nT. The static heading error is independent of the rotating field.
Table 3: The measurement results and the predictions are based on the rotating matrix simulation and density matrix simulation (static heading error).
@@ 2.3.1 Results of the systematics
Static heading error $B_{SH}$ 2.7 nT 2.7 nT
@@
Figure 2: (a) The measurement of the probe beam heading error. The solid points are experimental results, and the dashed lines are the theoretic plots based on our proposed model. (b) The magnetic field in the transverse direction is caused by the eddy current. The magnetic fields measured by both the Sine and Cosine switches are subtracted by 16.6 nT to remove the residual magnetic field.(c) The measurement of the eddy current induced magnetic field due to different switches. (d) The measurement of the eddy current time constant.
@@
[Dynamic heading error] In addition to the static heading error, a rotating magnetic field can induce a dynamic heading error. This effect is characterized in Table 4 using experimental and density matrix simulation results [34]. The density matrix model describes an ensemble of spins in a mixed state and is detailed in Secs. VI and VII of the supplemental material.
@@
When $1/T_{2}<\omega_{m}\ll\omega_{0}$, the dynamic heading error can be explained because the spin precession plane can adiabatically follow the rotation of the total magnetic field $\mathbf{B_{tot}}$. This keeps the relative angle $\Theta$ (as depicted in Fig. 2(b)) between $\mathbf{B_{tot}}$ and the spin precession plane constant, resulting in a constant precession frequency due to the weighted sum of different Zeeman transitions stays consistent. Thus, the dynamic heading error $B_{DH}=\mathcal{B}_{H}\cos\Theta$ depends on the initial angle $\Theta$ between the spin polarization and $\mathbf{B_{tot}}$, which remains constant and equal to its initial value.
Immediately after laser pumping, the initial spin $\mathbf{S}$ is along the negative y direction. $\mathbf{S_{\parallel}}$ and $\mathbf{S_{\perp}}$ are the spin projections of $\mathbf{S}$ along and perpendicular to the total magnetic field $\mathbf{B_{tot}}$ immediately after optical pumping, respectively.
If the $\mathbf{B_{tot}}$ starts initially perpendicular to $\mathbf{S}$ ( $\mathbf{S_{\parallel}}=0$), such as when $\mathbf{B_{m}}$ starts from the x-axis and $\mathbf{B_{tot}}$ is in the x-z plane, $\mathbf{S}$ is initially perpendicular to $\mathbf{B_{tot}}$ ($\Theta=\pi/2$), resulting in a dynamic heading error of 0. The spin precession plane can adiabatically follow the rotation of the total magnetic field, remaining perpendicular to it.
@@
However, if the $\mathbf{B_{tot}}$ initially starts non-perpendicularly to $\mathbf{S}$ ($\mathbf{S_{\parallel}}\neq 0$), such as when $\mathbf{B_{m}}$ starts from the negative y-axis and $\mathbf{B_{tot}}$ is initially in the y-z plane, $\mathbf{S}$ is no longer perpendicular to $\mathbf{B_{tot}}$, and there is an angle of $\Theta=\pi/2-\theta\approx 69^{\circ}$. This angle leads to a measured dynamic heading error of 2.6 nT, which is close to $\mathcal{B}_{H}\cos\Theta$. Similarly, if the rotating field starts from the positive y-axis ($\Theta=\pi/2+\theta$), the dynamic heading error becomes -2.6 nT. The full sign dependence of the dynamic heading error can be found in Table 2.
@@
Table 4: Measured dynamic heading error along the total magnetic field with $\beta\approx 0^{\circ}$, $\theta\approx 21^{\circ}$. “Exp” is the experimental data, “DM” is the density matrix simulation result.
$\mathbf{B_{m}}$ starts $\parallel\mathbf{S}$ $\mathbf{B_{m}}$ starts $\perp\mathbf{S}$
@@
The experimental results for the dynamic heading error are smaller than the density matrix simulation because the actual spin polarization is approximately 22% smaller than the simulation, which assumes near 100% polarization. For more data with different angles $\beta$ between the pump beam and the magnetic field, please refer to Sec. VII of the supplemental material.
@@
We also evaluated the dynamic heading error effect in the transverse directions. The summary of the experimental results and the theoretical predictions related to $\mathbf{B_{tot}}$ based on the density matrix are shown in Table 4. The dynamic heading error effect is close to zero in $B_{y}$ regardless of whether the rotating field starts parallel or perpendicular to $\mathbf{S}$. These small magnetic field variations may be caused by systematic mismatches in the threshold voltage of the MDA.
@@
The variation in $B_{x}$ measurement from the density matrix simulation in each shot is attributed to the probe beam heading error ($\tau_{\text{prob}}$). For instance, if the sensor has a misalignment such that $\alpha=1^{\circ}$, it leads to an approximate 1.2 nT probe beam heading error in the x-direction. However, these systematic errors can be eliminated through 4-shot averaging.
Additional data in Sec. VII of the supplemental material show that these magnetic field offsets do not change their signs when the polarization direction is reversed. This proves that they are not caused by the dynamic heading error.
@@
Overall, the rotating magnetic field can induce dynamic heading errors in total magnetic field measurements. These errors can be eliminated by either choosing the right starting phases for the rotating field, which starts the field initially perpendicular to the spin polarization, or by averaging the results of four shots, as shown in Table 2. The dynamic heading error does not contribute to the transverse magnetic fields, regardless of the starting phases of the rotating field.
@@
[Eddy current] The fast-altering rotating magnetic field can generate an eddy current in the electrically conductive magnetic shield. In our four-shot scheme that alternates the rotation direction, there are two switches in four shots, as shown in Fig. 1b. We discovered that how the rotating magnetic field is altered can affect the eddy current and its induced magnetic field. When the direction alteration of the rotating field happens at the peak values of the rotating magnetic field, from negative maximum to positive maximum and vice versa, we named this as the “cosine switch” (Fig. 1b top left). The average magnetic field caused by the eddy current equals zero. However, if the direction alteration of the rotating field happens at zero rotating magnetic field, we name it the “sine switch” (Fig. 1b top right), which would lead to a significant magnetic field caused by the eddy current.
@@
For instance, in y-direction, in Fig. 1b, the switch of $B_{y}$ occurs during the preparation time of the second shot, and another switch occurs during the preparation time of the fourth shot. Therefore, in the “sine switch” case, the second and fourth shots have the maximum eddy current, while the first and third shots have smaller eddy currents because they measure the decaying eddy current generated during the fourth and second shots. The sign of the magnetic field caused by the eddy current depends on the rotation direction of the rotating field. One panorama (80 shots) measurement result from the MDA is shown in Fig. 2(c). The sine switch can cause an eddy current magnetic field of approximately 450 nT, while the cosine switch can highly suppress the systematic of the eddy current. The small eddy current magnetic field in the cosine switch is due to the imperfection of the cosine switch. The instruments have a certain rising time for a sudden signal transition.
@@
The time constant of the eddy current is measured by applying a continuous rotating magnetic field with a constant phase (no switch) during one panorama measurement. As shown in Fig. 2(d), there are 79 data points representing 79 shot measurements. The first shot is dumped due to it overloading the phase shift range of the MDA. For each shot, it takes 1/120 s. The 60 Hz line frequency causes periodic variations in the magnified picture and its harmonics, which can be averaged out by one panorama measurement. We measured the time constant of the eddy current effect for our setup to be approximately 10.4 ms, with the corresponding cut-off frequency at 15 Hz.
@@
Therefore, we specifically developed a waveform as shown in Fig. 1b to ensure the rotation direction alteration always happens at the peak values of the adjacent rotating magnetic field (cosine switch). We can mostly eliminate the systematic effect caused by the eddy current by using the ”cosine switch” and 4-shot averaging.
2.3.2 Vector magnetic gradiometer
To demonstrate the magnetic gradiometer measurement of the FRF vector magnetometer, a multipass 87Rb cell is used. A high-power QPC laser polarizes the spins along the z-axis to increase spin the polarization, and $\pi/2$ magnetic field pulses are generated right after the pump to tip the spins from the z-axis to the y-axis. Two Carmel NK732 cards replace the MDA to acquire the rising and falling edge zero-crossings for 1 minute. The gradiometer sensitivity is shown in Fig. 3, with a sensitivity of 50 $\mathrm{fT/\sqrt{Hz}}$ for measuring the total field and sensitivities of 400 $\mathrm{fT/\sqrt{Hz}}$ for measuring $B_{x}$ and $B_{y}$, respectively. Their noise spectrums are flat down to 1 Hz and 0.1 Hz, respectively. For comparison, a scalar sensitivity of 40 $\mathrm{fT/\sqrt{Hz}}$ was achieved without applying any rotating field.

### v3

@@ II.3 Experimental results
The alkali cell is pumped by a sequence of pulses from a grating-stabilized diode laser, with a transverse relaxation time measured at $T_{2}\approx 3$ ms. Inside the magnetic shield, a set of three coils provides a rotating magnetic field and a leading magnetic field. The cell is heated to 100C by an electric heater driven by AC at a frequency of 131.5 kHz. The AC heater is turned on during the pump time and turned off during the measurement time to reduce magnetic noise from the heater.
As shown in Fig. 1a, a rotating magnetic field with an amplitude of approximately 18 $\mathrm{\mu T}$ and a freuqency of $\omega_{m}=2\pi\times 480$ Hz is applied in the transverse plane, and a leading magnetic field $B_{z}$ is applied along the longitudinal direction. The total amplitude of the applied magnetic field is maintained at approximately 50 $\mathrm{\mu T}$. The frequency of the pump beam pulse is set to 348 kHz, synchronized with the Larmor frequency to achieve the highest pumping rate. The duty cycle of a pump beam pulse is approximately 1.4%.
After the pump phase, the pump beam is switched off, and the free induction decay (FID) signal of the spin precession is measured using a linearly polarized probe beam. The probe beam, with a cross-sectional area of approximately 4 mm2 and a power of around 2 mW, is generated by a vertical cavity surface-emitting laser (VCSEL). The optical rotation of the linearly polarized probe beam is measured by a balanced polarimeter consisting of a Wollaston prism and a quadrant photodiode. The signal from the photodiode is then passed through a differential low-noise amplifier and a high-pass filter with a cutoff frequency of 150 kHz.
@@
II.3.1 Results of the systematics
The experimental results for the systematics are presented in Table 2.
Table 2: The measurement results and predictions are based on rotating matrix simulations and analytical models.
@@ II.3.1 Results of the systematics
Static heading error $B_{SH}$ ($\beta\approx 24^{\circ}$) 2.7 nT 2.9 nT
$|B_{PD}|$ ($\theta\approx 21^{\circ}$, $\beta\approx 0$) 2.6 nT 2.8 nT
[Berry’s phase] The Berry phase will result in a phase shift that adds to the total magnetic field measurement. When $\theta\approx 21^{\circ}$, magnetic field measurement systematic caused by Berry’s phase shift is measured to be 5.1 nT, while Eq. 6 predicts 4.7 nT ($\phi_{B}/\gamma$), as shown in Table. 2.
[Second harmonic phase shift] The amplitude of the second harmonic of $\omega_{m}$ in phase shift is 35 mrad, which matches Eq. 7 prediction well.
[Heading Errors] The static component of the dynamic heading error is measured when the magnetometer sensor rotates by an angle $\beta=24^{\circ}$ in the yz-plane, resulting in a heading error of $B_{SH}=\mathcal{B_{H}}\cos\theta\sin\beta$. From Eq. (8), the static heading error is estimated to be 2.9 nT, with an experimental result of 2.7 nT. The slightly lower experimental value is attributed to incomplete spin polarization.
@@
Figure 2: (a) The measurement of the probe beam heading error. The solid points are experimental results, and the dashed lines are the theoretic plots based on our proposed model. (b) The magnetic field in the transverse direction is caused by the eddy current. The magnetic fields measured by both the Zero-Altering and Peak-Altering are subtracted by 16.6 nT to remove the residual magnetic field.(c) The measurement of the eddy current induced magnetic field due to different switches. (d) The measurement of the eddy current time constant.
@@
A rotating magnetic field can induce a rotating-field-phase-dependent heading error, $B_{PD}$. This effect is characterized in Table 3 with both experimental results and density matrix simulations [32]. The density matrix model, which represents an ensemble of spins in a mixed state, is described in detail in Secs. VIII and IX of the supplemental material.
@@
When $1/T_{2}<\omega_{m}\ll\omega_{0}$, the dynamic heading error can be interpreted as the spin precession plane adiabatically following the rotation of the total magnetic field $\mathbf{B_{tot}}$. This keeps the relative angle between $\mathbf{B_{tot}}$ and the spin precession plane constant, resulting in a steady precession frequency. Consequently, the dynamic heading error depends on the initial angle between the spin polarization and $\mathbf{B_{tot}}$, which remains equal to its initial value. We define $\Theta$ as the angle by which $\mathbf{B_{tot}}$ deviates from the position where $\mathbf{S}$ is perpendicular to $\mathbf{B_{tot}}$, as shown in Fig. 2(b).
For example, when $\mathbf{B_{m}}$ begins along the x-axis, the spins are polarized in a direction tilted by an angle $\beta$ away from the negative y-axis in the yz-plane, and $\mathbf{B_{tot}}$ initially lies in the xz-plane. Due to symmetry, whether $\mathbf{B_{m}}$ starts from the positive or negative x-axis, the relative angles between $\mathbf{S}$ and $\mathbf{B_{tot}}$ remain the same. In this configuration, the dynamic heading error equals the static heading error, $\mathcal{B_{H}}\cos\theta\sin\beta$, with $B_{PD}=0$. Under these conditions, the spin precession plane follows the rotation of the total magnetic field while keeping $\Theta$ constant. Specifically, if $\beta=0$ initially, $\mathbf{S}$ stays perpendicular to $\mathbf{B_{tot}}$ throughout, resulting in $B_{DH}=0$.
@@
However, if $\mathbf{B_{m}}$ originates from the negative y-axis and $\mathbf{B_{tot}}$ is initially in the yz-plane, then $\mathbf{S}$ deviates from perpendicularity with $\mathbf{B_{tot}}$ by an angle $\Theta=\beta+\theta$. This results in a dynamic heading error of $\mathcal{B_{H}}\sin(\beta+\theta)$, with $B_{SH}=\mathcal{B_{H}}\cos\theta\sin\beta\approx 2.9$ nT, and $B_{DH}=\mathcal{B_{H}}\cos\beta\sin\theta\approx 2.5~{}\mathrm{nT}$. Conversely, if the rotating field begins from the positive y-axis, $\Theta=\beta-\theta$, with $B_{SH}$ remaining the same, but $B_{DH}=-\mathcal{B_{H}}\cos\beta\sin\theta\approx-2.5~{}\mathrm{nT}$. Table 1 provides a full overview of the sign dependence of the dynamic heading error.
@@
Table 3: Measured dynamic heading error along the total magnetic field with $\beta\approx 0^{\circ}$, $\theta\approx 21^{\circ}$. Here the static part of dynamic heading error $B_{SH}=0$. “Exp” is the experimental data, “DM” is the density matrix simulation result.
$\mathbf{B_{m}}$ initiates from Y-axis $\mathbf{B_{m}}$ initiates from X-axis ($\mathbf{B_{tot}}\perp\mathbf{S}$)
@@
The experimental results for the dynamic heading error are smaller than the analytical model and density matrix simulation because the actual spin polarization is smaller than the simulation, which assumes full polarization. For more data with different angles $\beta$ between the pump beam and the magnetic field, please refer to Sec. IX of the supplemental material.
@@
We further evaluated the dynamic heading error effect in the transverse directions. A summary of the experimental results and the theoretical predictions related to $\mathbf{B_{tot}}$, based on the density matrix, is provided in Table 3. The dynamic heading error effect remains close to zero in $B_{y}$, irrespective of whether the rotating field begins parallel or perpendicular to $\mathbf{S}$. These minor magnetic field variations may result from systematic mismatches in the MDA’s threshold voltage, leading to curvature that affects the fitting of the slope $\Phi_{s}$.
@@
The variation in $B_{x}$ measurement from the density matrix simulation in each shot is due to the probe beam heading error ($\phi_{\text{prob}}/\gamma$). For example, if the sensor is misaligned with an angle of $\alpha=1^{\circ}$, this results in an approximate 1.2 nT probe beam heading error in the x-direction. Additional data in Sec. IX of the supplemental material indicate that these magnetic field offsets maintain their signs even when the polarization direction is reversed, confirming that they are not caused by dynamic heading error. These systematic errors, however, can be eliminated through 4-shot averaging.
@@
Overall, the rotating magnetic field can induce dynamic heading errors in total magnetic field measurements, consisting of a static heading error component and a rotating-field-phase-dependent component. The rotating-field-phase-dependent heading error, $B_{PD}$, can be eliminated either by selecting appropriate starting phases for the rotating field—ensuring the field initially aligns perpendicularly or maintains a constant angle relative to the spins with alternating start directions—or by averaging the results of four shots, as shown in Table 1. Importantly, the dynamic heading error does not affect measurement of the transverse magnetic fields, regardless of the starting phases of the rotating field.
To investigate the probe beam heading error, the sensor rotates about the y-axis within the xz-plane at an angle of $\alpha$. This error can introduce a systematic effect in transverse magnetic field measurements that depends on the angle $\alpha$, with its sign determined by the rotation direction of the rotating field. The measured probe heading error at $\omega_{m}=2\pi\times 480$ Hz is shown in Fig.2(a), with the theoretical basis provided in Eq.9. “Theory 1” corresponds to anti-clockwise rotations (Shots 1 and 3 in Table 1), given by $-\omega_{m}/\gamma\cdot\tan\alpha$, while “Theory 2” represents clockwise rotations (Shots 2 and 4 in Table 1), given by $\omega_{m}/\gamma\cdot\tan\alpha$. The probe beam heading error can be effectively canceled by averaging the measurements from two opposite rotating field directions.
@@
[Eddy current] The alternating fast rotating magnetic field can generate an eddy current in the electrically conductive magnetic shield. In our four-shot scheme, which alternates the rotation direction, there are two switches within four shots, as shown in Fig.1b. We found that the way the rotating magnetic field is altered affects the eddy current and its induced magnetic field. When the direction of the rotating field changes at the peak values of the field—from negative maximum to positive maximum and vice versa—we refer to this as “Peak-Altering” (Fig.1b, top left). In this case, the average magnetic field caused by the eddy current equals zero. However, if the direction change occurs when the rotating magnetic field is at zero, termed “Zero-Altering” (Fig. 1b, top right), it leads to a significant magnetic field caused by the eddy current.
@@
In our setup, the y-direction switch of $B_{y}$ occurs during the preparation times of the second and fourth shots, as shown in Fig.1b. In the “Zero-Altering” case, the second and fourth shots experience the maximum eddy current, while the first and third shots have smaller eddy currents, as they capture the decaying eddy current generated during the fourth and second shots. The sign of the eddy current-induced magnetic field depends on the rotation direction of the rotating field. A panorama measurement (80 shots) result from the MDA, shown in Fig.2(c), indicates that Zero-Altering can cause an eddy current magnetic field of approximately 450 nT. In contrast, Peak-Altering effectively suppresses the systematic effect of the eddy current. Any remaining small eddy current magnetic field in Peak-Altering is due to imperfections, as instruments have a finite rise time during abrupt signal transitions.
@@
To characterize the time constant of the eddy current, we applied a continuous rotating magnetic field with a constant phase (no alteration) over one panorama measurement. In Fig. 2(d), 79 data points represent 79 individual shot measurements, with the first shot discarded due to phase shift overload in the MDA. Each shot measurement lasts 1/120 s. The 60 Hz line frequency produces periodic variations and harmonics, which are effectively averaged out over one panorama. Our setup’s measured eddy current time constant is approximately 10.4 ms, corresponding to a cut-off frequency of 15 Hz.
@@
To mitigate the eddy current’s systematic effect, we designed a waveform shown in Fig. 1b to ensure that rotation direction alterations consistently occur at the peak values of the rotating magnetic field (Peak-Altering). By employing Peak-Altering and 4-shot averaging, we effectively reduce the eddy current’s systematic effect.
II.3.2 Results of magnetic field sensitivities
A magnetic field gradiometer measures the variation in magnetic field strength over a short distance, rather than the absolute field strength at a single location. This method is effective for detecting local magnetic changes while reducing the influence of uniform, distant magnetic interference [39].To demonstrate the gradiometric measurement capability of the FRF vector magnetometer, we configured a setup using a multipass 87Rb cell and a high-power QPC laser with a 25 W output. The laser, operating in pulsed mode, polarizes the spins along the z-axis to enhance spin alignment and is then turned off during the readout phase. Following this pumping stage, $\pi/2$ magnetic field pulses are applied to tip the spins from the z-axis to the y-axis.
Data acquisition is performed with two Carmel NK732 cards in place of the MDA, recording the zero-crossing data—including both frequency and phase information—from two magnetometers over a one-minute interval. We derive the gradiometer signal by subtracting the zero-crossing data from the two magnetometers, forming a first-order gradiometer that cancels common-mode magnetic field noise. Assuming that any remaining noise is uncorrelated, we further divide this residual noise by $\sqrt{2}$ to assess the intrinsic magnetic field sensitivity of each channel [39].
The gradiometer sensitivities are shown in Fig. 3, with a sensitivity of 35 $\mathrm{fT/\sqrt{Hz}}$ for the total field and 280 $\mathrm{fT/\sqrt{Hz}}$ for the $B_{x}$ and $B_{y}$ components, respectively. Their noise spectrums are flat down to 1 Hz and 0.1 Hz, respectively. For comparison, a scalar sensitivity of 28 $\mathrm{fT/\sqrt{Hz}}$ was achieved without the application of a rotating field.